OpenStudy (steve816):

Ryan has a savings account that pays 4.5% interest annually compounded quarterly. She has not made any deposits or withdrawals for many years. Suppose that her current balance is $4,516.32. How much money was in her account 5 years ago?

OpenStudy (arindameducationusc):

Use compound interest formula

OpenStudy (steve816):

I always forget that formula xD

OpenStudy (steve816):

someone walk me through please!

OpenStudy (dan815):

this is why i think econ is so bs okay.. they say 4.5% annual interest compounded quarterly that in no way tells you that it means that every quarter the compound interest is 4.5/4 %

OpenStudy (arindameducationusc):

@steve816 you there?

OpenStudy (dan815):

now in 5 years that means the inital money was compounded 5*4 =20 times of 4.5/4%

OpenStudy (dan815):

let x be the money in the beginning x*(1+4.5/4%)^20=x*(1+0.045/4)^20=4,516.32 x*(1+0.045/4)^20=4,516.32 x=4,516.32/(1+0.045/4)^20

OpenStudy (dan815):

you can see why this formula comes about if you try do compound x one at a time (x*(1+r)) money after 1 compound (x*(1+r))*(1+r) = x*(1+r)^2 money after 2nd compound (x*(1+r))*(1+r)*(1+r)=x*(1+r)^3 money after 3rd compound and so on . . x*(1+r)^20 = money after 20 compounds